eBook The exercise involving data in this and subsequent sections were designed to be solved using Excel. The following estimated regression equation is based on 30 observations. ? 17.3 + 3.5c,-2x2 + 7.4x3 + 2.8z4 The values of SST and SSR are 1,801 and 1,755, respectively. a. Compute R2 (to 3 decimals) .026 b. Compute R (to 3 decimals). .129 c. How good is the fit provided by the estimated regression equation? The estimated regression equation provided an excellent fit Hide Feedback Partially Correct
Answer #1

a) R2 = SSR/SST = 1755/1801 = 0.974

b) Adjusted R2:

R_a^2 = 1 - \frac{(1-R^2)(N-1)}{N-p-1}

R_a^2 = 1 - \frac{(1-0.974)(30-1)}{30-4-1}

R_a^2 = 0.970

c) The model is a good fit.